Find the distance of the pt. (1,2) from the straight line with slope 5 and passing through the pt. Of intersection of x+2y=5 and x-3y=7.

Grade:11

2 Answers

Vitthal Gupta
6 years ago
Let the line beY=mx+cm is slope and c is a constantThen solve both the lines given to get point of intersection from which the unknown line is also passingNow satisfy the points in unknown line (y=mx+c)Then you get the equation of lineNow use formula to get distance(ax+by+c)/√(a^2+b^2)Where x,y is the coordinates of the given point
Priya
6 years ago
Line having slope=5 perpendicular to points 1,2 the equation of line becomes x+5y-11=0(m=-1/5 point 1,2) now solving x+5y-11=0 and 25x-5y-147=0 we get x=79/13 y=64/65 now find distance between (1,2) and(79/13,64/65) is the req distance